OpenStudy (anonymous):

If $240 is invested at an interest rate of 9% per year and is compounded monthly, how much will the investment be worth in 14 years? $68.39 $704.28 $842.13 $846.10

hero (hero):

Monthly Compound Interest Formula: \[A = P\left(1 + \frac{r}{12}\right)^{12t}\] A=Account Balance P=Principal r=rate (decimal form) t=time (in years)